OpenStudy (darkprince14):

Prove: If A and B are finite sets than A U B is a finite set and card(AUB) = card(A) + card(B)

OpenStudy (anonymous):

Suppose \(A=B\), is it true?

OpenStudy (anonymous):

\[ \|A\cup B\| = \|A\| + \|B\| - \|A\cap B\| \]

OpenStudy (darkprince14):

Nope, A is not equal to B

OpenStudy (darkprince14):

It just said that they're finite..

OpenStudy (anonymous):

Two finite sets can be equal. I'm telling you about a simple counter example.

OpenStudy (darkprince14):

Oooh... sorry for that...

OpenStudy (zzr0ck3r):

if they are both finite and nonempty, then there exists a bijection(f) \[f:(1,2....m)\rightarrow A\\and\\g:(1,2,....n)\rightarrow B\] now define a new fuction\[h:(1,2,....,m+n)\rightarrow A\cup B\\where\space h(i)=i,for\space i=1,2,...m\\and\\h(i)=g(i-m)\space for\space i=m+1,m+2,....m+n \] now show that h is surjective.

OpenStudy (akashdeepdeb):

Are you talking about there cardinal numbers?

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

hmm let A = {1},|A| = 1 let B = {1},|B| = 1 AUB = {1} but|A|+|B| = 2

OpenStudy (zzr0ck3r):

@darkprince14 you understand? I dont think you can prove that last part.

terenzreignz (terenzreignz):

Maybe this instead... \[\Large |A \cup B | \le |A| +|B|\]

OpenStudy (zzr0ck3r):

it must be

OpenStudy (darkprince14):

@satellite73 please help..

OpenStudy (zzr0ck3r):

what do you need help with?